OpenStudy (anonymous):

Precalc: Find the foci and vertices of the ellipse with equation: x^2/4+y^2/16=1

OpenStudy (dumbcow):

ok from general ellipse equation we know that center is (0,0) a^2 = 16 --> a = 4 b^2 = 4 --> b = 2 Find Foci a^2 = b^2 +c^2 16 = 4 + c^2 c^2 = 12 c = 2sqrt(3) Vertices: (0,-4) and (0,4) Foci: (0,-2sqrt3) and (0,2sqrt3)